# The Expressive Capacity of State Space Models: A Formal Language Perspective

### v2

@@ Abstract
Recently, recurrent models based on linear state space models (SSMs) have shown promising performance in language modeling (LM), competititve with transformers. However, there is little understanding of the in-principle abilities of such models, which could provide useful guidance to the search for better LM architectures. We present a comprehensive theoretical study of the capacity of such SSMs as it compares to that of transformers and traditional RNNs. We find that SSMs and transformers have overlapping but distinct strengths. In star-free state tracking, SSMs implement straightforward and exact solutions to problems that transformers struggle to represent exactly. They can also model bounded hierarchical structure with optimal memory even without simulating a stack. On the other hand, we identify a design choice in current SSMs that limits their expressive power. We discuss implications for SSM and LM research, and verify results empirically on a recent SSM, Mamba.
@@ 1 Introduction
After their introduction [69], transformers rapidly became the primary workhorse of NLP, powering most of today’s large language models (LLMs). Compared to previously-dominant recurrent architectures [RNNs 17, 29], transformers offered a key advantage: parallelized training by avoiding recurrence. However, building on a long history of continuous dynamical models [e.g. 34, 35] and early work on faster RNNs [8, 41], a recent line of work has developed state space models (SSMs) rivaling the performance of transformers [e.g. 24, 23, 67, 14, 72, 56]. These SSMs are recurrent models that—while formulated in terms of iterative state updates—allow efficient parallelization.
@@
One common approach to understanding the capabilities of computational architectures is through their expressive capacity in simulating automata and modeling language classes; indeed, a sizeable literature has studied transformers [e.g. 54, 25, 6, 73, 44, 45, 15, 66, 10, 59, 53] and RNNs [e.g. 62, 31, 32, 70, 28] through this lens. As the difficulty of many computational problems is well-understood in terms of such language classes, results about expressive capacity directly yield results about the ability to model specific computational problems.
While a substantial number of results have been obtained for transformers and traditional RNNs, understanding remains largely open for SSMs. In an initial step, Merrill et al. [49] showed that all problems computable by SSMs are contained in $\operatorname{TC}^{0}$, a circuit complexity class that is known to also cover transformers [48, 65]. Under standard conjectures, this suggests that certain types of state tracking are hard for both models. Jelassi et al. [33] provided evidence for differences between the architectures, showing that transformers are better than SSMs at the specific problem of copying strings – a problem well within $\operatorname{TC}^{0}$. However, beyond these results, broader detailed understanding of the power of SSMs and how they compare to RNNs and transformers remains open.
Our contribution in this paper is to provide rigorous understanding of SSMs’ abilities in different classes of languages. We show that transformers and SSMs cover overlapping but distinct fragments of $\operatorname{TC}^{0}$. For instance, SSMs can model bounded hierarchical structure in ways similar to transformers and traditional RNNs, even without embedding a stack-like structure (Theorem 6). For regular languages involving modular counting, such as the PARITY function (Theorem 2), we identify a design choice that makes extant SSMs struggle in ways similar to transformers. In other cases, we show that SSMs resolve a failure case of transformers: they effortlessly model Flip Flop state tracking (Theorem 1). We discuss take-aways for SSM and LLM research in Section 5; among others, our results suggest future LM architectures might need to combine both attention and state spaces.
2 Background: State Space Models
@@
[…]
@@
where $\operatorname{Mix}_{j}(\cdot)$ can contain linear or (Swi)GLU components [e.g. 56, 23]. We will take $\operatorname{Norm}$ to implement RMSNorm [75]; LayerNorm [4] can be covered by absorbing centering into $\operatorname{Mix}_{2}$.
@@ A Full SSM
In Mamba, (1) directly maps onto Eqs. (2a) and (2b) in Gu and Dao [23]. The notation of Gu and Dao [23] use a matrix multiplication $\overline{A}h_{t-1}$ instead of elementwise multiplication $A(x_{t})\circ h_{t-1}$ in (1), but importantly, Mamba’s $\overline{A}$ is diagonal, so we can take $A(x_{t})_{i}=\overline{A}_{ii}$. Some SSMs assume nondiagonal $A(x_{t})$, but typically this matrix is diagonalizable [e.g. 24, 67], so that the SSM is still equivalent to one of the form (1). We discuss how other SSMs instantiate (1) in Appendix A. Some models assume complex-valued activations (Appendix A); our results largely do not depend on this distinction, but take it into account where needed (Theorem 13). Some SSMs [e.g. 23] use different numbers of channels in $x_{t}$ and $h_{t}$ using state expansion; as this does not affect expressive capacity, we will simply assume a constant dimensionality $d$. Local convolutions [e.g. 21] can be simulated with an SSM layer and do not increase expressive capacity (Remark 18).
We will find that two design choices have nontrivial impact on expressive capacity: The first one is time invariance: we call an SSM time-invariant if $A(x_{t})$ does not depend on $x_{t}$. Some SSMs, such as S4 [24] and Retnet [67] are time-invariant; Mamba [23], Griffin [14], GLA [72], HGRN [57, 56], QRNN/SRU [8, 41] are not (Appendix A). The second one is the sign of the entries of $A(x_{t})$: Across all non-time-invariant SSMs surveyed, we find that the gate is always nonnegative (Appendix A): $A(x_{t})\geq 0$ (nonnegative) due to exponential or sigmoid parameterizations of the gate – this choice turns out to limit expressive capacity (Theorem 2).
@@
While the abstract form (12) is common across the SSM literature, differences in parameterization may have substantial effect on efficiency and training stability. Studying expressiveness allows us to abstract away from these differences to a remarkable degree: We will allow $A,B,\rho$ to be arbitrary functions with the given input-output properties. Our negative results are based on abstract properties of the setup (12), which fundamentally bottlenecks SSMs through elementwise linear state updates. For our positive results, will use empirical learnability experiments to verify that learnable solutions instantiating them (though not necessarily implementing the same constructions as used in the proofs) do exist in a recent SMM [Mamba, 23].
Traditional RNNs
We contrast SSMs with traditional RNNs such as simple RNNs or LSTMs: for these, the recurrence in Eq. (1) is replaced by
$h_{t}=\psi(h_{t-1},x_{t})$ (3)
where $\psi$ could be linear, an MLP [17], or a more complex gated function [29].
@@
While Eq.(1) assumes arbitrary real-valued activations, real-world implementations can only represent numbers with bounded precision. Formally, we adopt the finite precision notion used by Weiss et al. [70] in a study of the expressive power of traditional RNNs: We allow an unbounded number of integer bits, but only $p$ fractional bits, independent of the length of the input. See Appendix E for discussion.
@@ 3 Theoretical Results
3.1 Motivation
We study three foundational types of data structures needed for modeling formal languages [30]: finite state automata (Theorem 1, 2, 4), counters (Theorem 5), and stacks (Theorem 6). All of these data structures can be understood in two equivalent forms: One is to track a state sequence over an input, where each input symbol engenders a specific transformation on the state. The other one—more commonly considered in research on expressive capacity and closer to language modeling–considers formal languages—sets of finite strings that are defined by the property that an automaton reaches one of a pre-specified set of “accepting” states after traversing the word. We will focus on the latter perspective, enabling easy comparison with existing results on transformers and RNNs.
A finite-state-automaton (see Definition 7) represents a general state tracking problem over a finite state space, without imposing further structure on the state space: The automaton keeps track of a single state from a finite state space; when reading a string from left to right, each symbol engenders a specific transformation of the state. At each position, the current state determines which symbols can come next; membership in a formal language is determined by the state reached after reading the full string. Finite-state-automata are equivalent in expressivity to regular expressions, and define the regular languages [37]. Indeed, we will be able to provide a precise criterion identifying which such finite state tracking problems–or equivalently, regular languages–SSMs such as Mamba are capable of in the finite-precision setting (Theorem 4).
Allowing an automaton to keep track of one or more counters [20]—integers that are incremented or decremented at each symbol read—turn the state space infinite, but in a highly structured manner. SSMs can model this datastructure (Theorem 5), as can RNNs and transformers [70, 6]. Stacks, a first-in-first-out datastructure, enable automata to keep track of hierarchical structure, foundational to natural language [11]. We show that SSMs can implement shortcut solutions to bounded hierarchical structure even without implementing a stack (Theorem 6) – these are likely to be most useful to natural language given the boundedness of human memory [50, 36].
3.2 Background: Formal Language Prediction and Recognition
We next need to define what it means for an SSM to model a formal language. A standard notion is the notion of recognition, where the task is to classify a full string as belonging to the language or not. Formally, for an SSM with $d_{pred}=1$, we say that it recognizes a language $L$ if the output $\rho(z_{|{\bf w}|}^{(L)})$ equals—when the SSM is run on ${\bf w}\in\Sigma^{*}$—1 if ${\bf w}\in L$ and 0 else.
However, such a classification task is arguably not always matched to dominant use cases in predictive sequence modeling, where the task is to predict the next token at each step. Thus, we also cast formal languages into a language modeling and sequence prediction framework. We adopt the task of Bhattamishra et al. [6], where the model is asked to output at each step in a sequence the set of possible next symbols. Let $\operatorname{Prefix}(L):=\{w:w\in\Sigma^{*},w\Sigma^{*}\cap L\neq\emptyset\}$ the set of valid prefixes of $L$. We then say that a model predictively models a language $L$ if (Figure 1), given a valid prefix $w\in\operatorname{Prefix}(L)$, it outputs the finite set
$\{\sigma\in\Sigma:w\sigma\Sigma^{*}\cap L\neq\emptyset\}$ (4)
We think of each such set as an atomic label; the set of possible labels is the power set of the finite alphabet $\Sigma$. Here, $d_{pred}=2^{|\Sigma|}$. Predictive modeling can be easily converted into recognition by checking whether any symbol in the sequence is not in the predictive set at the preceding position; this can be done by adding 1 SSM layer. Conversely, if we can show that SSMs cannot recognize a language, this proves they also cannot perform predictive modeling for it, as they then cannot correctly predict where EOS can appear. To get the strongest results, we thus prove positive results for predictive modeling, and negative results for recognition.
@@ 3.3 Length-Generalizing Representations for Flip-Flop State Tracking
As motivated in Section 3.1, we begin with finite-state-automata – equivalently, regular languages, a well-studied setting for understanding the expressive power of transformers [e.g. 25, 6, 3, 44, 45] and a fundamental model of state tracking when the number of possible states is finite [45, 49]. Traditional RNNs can emulate all finite-state-automata at finite precision [31, 32]. Starting from two specific languages, Flip Flop (Section 1) and PARITY (Section 2), we derive an exact characterization of the regular languages modeled by a broad class of SSMs at finite precision (Theorem 4).
We begin on the positive side, by establishing a case where SSMs pattern with RNNs in avoiding a failure mode of self-attention. The Flip Flop languages [45] are a simple instance of state tracking defined in terms of write, read, and ignore instructions. Each write instruction comes with a piece of information; whenever a read instruction is encountered, the information written by the last write instruction is recalled. Formally, define $\mathcal{L}_{FF}$ to be the set of finite strings ${\bf x}$ over $\Sigma=\{{\tt r},{\tt w},{\tt i},0,1\}$, where $x_{1},x_{3},\dots\in\{{\tt r},{\tt w},{\tt i}\}$, $x_{2},x_{4},\dots\in\{0,1\}$, and where the bit following any ${\tt r}$ matches the bit following the last preceding occurrence of ${\tt w}$. Liu et al. [44] show that the Flip Flop language, as an abstraction, is a fundamental ingredient of many long-range reasoning settings. It can be represented with a small finite-state-automaton, and LSTMs learn $\mathcal{L}_{FF}$ well [45]. Transformers can in principle represent it [44, 45], though known constructions are not inherently length-generalizing. Indeed, transformers empirically struggle in generalization [45]; intuitively, this may happen because attention heads aggregate information in a commutative manner, and reliably attending to the last write instruction requires strong position dependence in the attention weights. SSMs pattern with traditional RNNs: they can easily represent Flip Flop at arbitrary input lengths:
@@
A two-layer SSM with finite precision predictively models $\mathcal{L}_{FF}$ at arbitrary lengths.
In the construction, the first layer records the last instruction token, achieved in (1) by setting $A(e({\tt r}))=A(e({\tt w}))=A(e({\tt i}))=0$, and $A(e(0)=A(e(1))=1$, and setting $B(e(0))=B(e(1))=0$. Additional dimensions forward the current token to $h_{t}^{(1)}$. In the output of the first layer $z_{t}^{(1)}$, whenever the input is 0 or 1, the model now has access both to the current token $w_{t}$ and the preceding token $w_{t-1}$, which must have been an instruction. Based on this information, the model can set the gate to overwrite the state $h_{t-1}^{(2)}$ with the current input token when the preceding token was ${\tt w}$, and pass along the state $h_{t-1}^{(2)}$ unaltered otherwise. This, together with $z_{t}^{(1)}$, is sufficient for always identifying the legal next symbols in $\mathcal{L}_{FF}$. The formal proof is in Appendix B.1.

### v3

@@ Abstract
Recently, recurrent models based on linear state space models (SSMs) have shown promising performance in language modeling (LM), competititve with transformers. However, there is little understanding of the in-principle abilities of such models, which could provide useful guidance to the search for better LM architectures. We present a comprehensive theoretical study of the capacity of such SSMs as it compares to that of transformers and traditional RNNs. We find that SSMs and transformers have overlapping but distinct strengths. In star-free state tracking, SSMs implement length-generalizing solutions to problems that transformers struggle to represent exactly. They can also model bounded hierarchical structure with optimal memory even without simulating a stack. On the other hand, we identify a design choice in current SSMs that limits their expressive power. We discuss implications for SSM and LM research, and verify results empirically  [^1] Code is available at: https://github.com/lacoco-lab/ssm_expressivity on a recent SSM, Mamba.
@@ 1 Introduction
Transformers (vaswani2017attention) power most large language models (LLMs) today, as they offer the advantage of parallelized training by avoiding recurrence, compared to the previously dominant recurrent achitectures (RNNs elman1990finding; hochreiter1997long). However, building on a long history of continuous dynamical models (e.g. kalman1960general; kalman1963mathematical) and work on faster RNNs (bradbury2016quasi; lei2017simple), a recent line of work has developed state space models (SSMs) rivaling the performance of transformers (e.g. gu2021efficiently; Gu2023Mamba; sun2023retentive; De2024Griffin; yang2023gated; qin2024hgrn2). These SSMs are recurrent models, formulated in terms of iterative state updates, while still allowing efficient parallelization.
@@
One common approach to understanding the capabilities of computational architectures is through their expressive capacity in simulating automata and modeling language classes; indeed, a sizeable literature has studied transformers (e.g. perezturing; hahn2020theoretical; bhattamishra2020ability; Yao_2021; liu2022transformers; Liu2023FlipFlop; deletang2022neural; strobl2023survey; chiang2023tighter; sanford2024representational; peng2024limitations) and RNNs (e.g. siegelman1991neural; horne1993bounds; indyk1995optimal; weiss2018practical; hewitt2020rnns) through this lens. As the difficulty of many computational problems is well-understood in terms of such language classes, results about expressive capacity directly yield results about the ability to model specific computational problems.
While a substantial number of results have been obtained for transformers and traditional RNNs, understanding remains largely open for SSMs. In an initial step, merrill2024illusion showed that all problems computable by SSMs are contained in $\operatorname{TC}^{0}$, a circuit complexity class that is known to also cover transformers (merrill2023logic; strobl2023averagehard). Under standard conjectures, this suggests that certain types of state tracking are hard for both models. jelassi2024repeat and bhattamishra2024separations provided evidence of differences between these architectures, showing that transformers outperform SSMs on copying or retrieving from long strings–tasks well within $\operatorname{TC}^{0}$. zubic2024limits showed that multi-layer SSMs are constrained by their logarithmic space computational capacity, limiting their ability at algorithmic tasks such as multi-digit multiplication.
However, a more fine-grained understanding of the power of SSMs, and how they compare to RNNs and transformers, remains an open question. Our contribution in this paper is to provide rigorous understanding of SSMs’ abilities in different classes of languages. We show that transformers and SSMs cover overlapping but distinct fragments of $\operatorname{TC}^{0}$. For instance, SSMs can model bounded hierarchical structure in ways similar to transformers and traditional RNNs, even without embedding a stack-like structure (Theorem 6). For regular languages involving modular counting, such as the PARITY function (Theorem 2), we identify a design choice that makes extant SSMs struggle in ways similar to transformers. In other cases, we show that SSMs resolve a failure case of transformers: they effortlessly model Flip Flop state tracking (Theorem 1). We discuss take-aways for SSM and LLM research in Section 5; among others, our results suggest future LM architectures might need to combine both attention and state spaces.
2 Background
2.1 State Space Models
@@
$h_{0}\in$ $\mathbb{R}^{d}$ $B(x_{t})\in\mathbb{R}^{d}\ \text{(increment)}$
$A(x_{t})\in$ $\mathbb{R}^{d}\ \text{(gate)}$ $\phi:\mathbb{R}^{2d}\rightarrow\mathbb{R}^{d}\ \text{(transform)}$
@@
where $\operatorname{Mix}_{j}(\cdot)$ can contain linear or (Swi)GLU components (e.g. qin2024hgrn2; Gu2023Mamba). We will take $\operatorname{Norm}$ to implement RMSNorm zhang2019root; LayerNorm ba2016layer can be covered by absorbing centering into $\operatorname{Mix}_{2}$.
@@ A Full SSM
In Mamba, (1) directly maps onto Eqs. (2a) and (2b) in Gu2023Mamba. The notation of Gu2023Mamba use a matrix multiplication $\overline{A}h_{t-1}$ instead of elementwise multiplication $A(x_{t})\circ h_{t-1}$ in (1), but importantly, Mamba’s $\overline{A}$ is diagonal, so we can take $A(x_{t})_{i}=\overline{A}_{ii}$. Some SSMs assume nondiagonal $A(x_{t})$, but typically this matrix is diagonalizable (e.g. gu2021efficiently; sun2023retentive), so that the SSM is still equivalent to one of the form (1). We discuss how other SSMs instantiate (1) in Appendix A. Some models assume complex-valued activations (Appendix A); our results largely do not depend on this distinction, but take it into account where needed (Theorem 13). Some SSMs (e.g. Gu2023Mamba) use different numbers of channels in $x_{t}$ and $h_{t}$ using state expansion; as this does not affect expressive capacity, we will simply assume a constant dimensionality $d$. Local convolutions (e.g. fu2023hungry) can be simulated with an SSM layer and do not increase expressive capacity (Remark 19).
We will find that two design choices have nontrivial impact on expressive capacity: The first one is time invariance: we call an SSM time-invariant if $A(x_{t})$ does not depend on $x_{t}$. Some SSMs, such as S4 (gu2021efficiently) and Retnet (sun2023retentive) are time-invariant; Mamba (Gu2023Mamba), Griffin (De2024Griffin), GLA (yang2023gated), HGRN (qin2024hierarchically; qin2024hgrn2), QRNN/SRU bradbury2016quasi; lei2017simple are not (Appendix A). The second one is the sign of the entries of $A(x_{t})$: Across all non-time-invariant SSMs surveyed, we find that the gate is always nonnegative (Appendix A): $A(x_{t})\geq 0$ (nonnegative) due to exponential or sigmoid parameterizations of the gate – this choice turns out to limit expressive capacity (Theorem 2).
@@
While the abstract form (12) is common across the SSM literature, differences in parameterization may have substantial effect on efficiency and training stability. In particular, the parameterization of $A(x_{t})$ has been the subject of substantial research (e.g. gu2020hippo; gu2021efficiently; yu2023robustifying; wang2023stablessm). However, studying expressiveness allows us to abstract away from these differences to a remarkable degree: We will allow $A,B,\rho$ to be arbitrary functions with the given input-output properties. Our negative results are based on abstract properties of the setup (12), which fundamentally bottlenecks SSMs through elementwise linear state updates. For our positive results, will use empirical learnability experiments to verify that learnable solutions instantiating them (though not necessarily implementing the same constructions as used in the proofs) do exist in a recent SSM (Mamba, Gu2023Mamba).
We contrast SSMs with traditional RNNs such as simple RNNs or LSTMs: for these, the recurrence in Eq. (1) is replaced by $h_{t}=\psi(h_{t-1},x_{t})$ where $\psi$ could be linear, an MLP (elman1990finding), or a more complex gated function (hochreiter1997long).
@@
While Eq.(1) assumes arbitrary real-valued activations, real-world implementations can only represent numbers with bounded precision. Formally, we adopt the finite precision notion used by weiss2018practical in a study of the expressive power of traditional RNNs: We allow an unbounded number of integer bits, but only $p$ fractional bits, independent of the length of the input. See Appendix E for discussion.
2.2 Modeling Formal languages
We study three foundational types of data structures needed for modeling formal languages (hopcroft2001introduction): finite state automata (Theorem 1, 2, 4), counters (Theorem 5), and stacks (Theorem 6). These data structures can be understood in two equivalent forms: One is to track a state sequence over an input, where each input symbol engenders a specific transformation on the state. The other one, more commonly considered in research on expressive capacity, considers formal languages—sets of finite strings that are defined by the property that an automaton reaches one of a pre-specified set of “accepting” states after traversing the word. We focus on the latter, enabling easy comparison with existing results on transformers and RNNs.
A finite-state-automaton (see Definition 7) represents a general state tracking problem over a finite state space, without imposing further structure on the state space: The automaton keeps track of a single state from a finite state space; when reading a string from left to right, each symbol engenders a specific transformation of the state. At each position, the current state determines which symbols can come next; membership in a formal language is determined by the state reached after reading the full string. Finite-state-automata are equivalent in expressivity to regular expressions, and define the regular languages (kleene1951representationof).
Allowing an automaton to keep track of one or more counters (Fischer1968Counter)—integers that are incremented or decremented at each symbol read—turns the state space infinite, but in a highly structured manner. SSMs can model this datastructure (Theorem 5), as can RNNs and transformers (weiss2018practical; bhattamishra2020ability). Stacks, a first-in-first-out datastructure, enable automata to keep track of hierarchical structure, foundational to natural language (lees1957syntactic). We show that SSMs can implement shortcut solutions to bounded hierarchical structure even without implementing a stack (Theorem 6) – these are likely to be most useful to natural language given the boundedness of human memory (miller1963finitary; karlsson2007constraints).
2.3 Formal Language Prediction and Recognition
We next need to define what it means for an SSM to model a formal language. The notion of recognition, where the task is to classify a full string as belonging to the language or not. Formally, for an SSM with $d_{pred}=1$, we say that it recognizes a language $L$ if the output $\rho(z_{|{\bf w}|}^{(L)})$ equals—when the SSM is run on ${\bf w}\in\Sigma^{*}$—1 if ${\bf w}\in L$ and 0 else.
However, such a classification task is arguably not always matched to dominant use cases in predictive sequence modeling, where the task is to predict the next token at each step. Thus, we also cast formal languages into a language modeling and sequence prediction framework. We adopt the task of bhattamishra2020ability, where the model is asked to output at each step in a sequence the set of possible next symbols. Let $\operatorname{Prefix}(L):=\{w:w\in\Sigma^{*},w\Sigma^{*}\cap L\neq\emptyset\}$ the set of valid prefixes of $L$. We then say that a model predictively models a language $L$ if (Figure 1), given a valid prefix $w\in\operatorname{Prefix}(L)$, it outputs the finite set
$\{\sigma\in\Sigma:w\sigma\Sigma^{*}\cap L\neq\emptyset\}$ (3)
We think of each such set as an atomic label; the set of possible labels is the power set of the finite alphabet $\Sigma$ (here, $d_{pred}=2^{|\Sigma|}$). Importantly, in both recognition and predictive modeling, we test the SSMs’ ability across arbitrary input lengths, i.e. the choice of input length does not affect the inherent capability to recognize or predictively model the language. Predictive modeling can be easily converted into recognition by checking whether any symbol in the sequence is not in the predictive set at the preceding position; this can be done by adding 1 SSM layer. Conversely, if we can show that SSMs cannot recognize a language, this proves they also cannot perform predictive modeling for it, as they then cannot correctly predict where EOS can appear. To get the strongest results, we thus prove positive results for predictive modeling, and negative results for recognition.
@@ 3.1 Length-Generalizing Representations for Flip-Flop State Tracking
Flip Flop languages (Liu2023FlipFlop) are a simple instance of state tracking defined in terms of write, read, and ignore instructions. Each write instruction comes with a piece of information; whenever a read instruction is encountered, the information written by the last write instruction is recalled. Formally, $\mathcal{L}_{FF}$ is the set of finite strings $\bf x$} over $\Sigma=\{{\tt r},{\tt w},{\tt i},0,1\}$, where $x_{1},x_{3},\dots\in\{{\tt r},{\tt w},{\tt i}\}$, $x_{2},x_{4},\dots\in\{0,1\}$, and where the bit following any $\tt r$} matches the bit following the last preceding occurrence of $\tt w$}. liu2022transformers show that the Flip Flop language, as an abstraction, is a fundamental ingredient of many long-range reasoning settings. It can be represented with a small finite-state-automaton, and LSTMs learn $\mathcal{L}_{FF}$ well (Liu2023FlipFlop). Transformers can in principle represent it (liu2022transformers; Liu2023FlipFlop), though known constructions are not inherently length-generalizing, a fact confirmed empirically; intuitively, this may happen because attention heads aggregate information in a commutative manner, and reliably attending to the last write instruction requires strong position dependence in the attention weights. SSMs, similar to traditional RNNs can easily represent Flip Flop at arbitrary input lengths and thus avoid a failure mode of self attention:
@@
There is a two-layer SSM that predictively models $\mathcal{L}_{FF}$ at all lengths, at finite precision.
In the construction (Figure 3), the first layer records the last instruction token, achieved in (1) by setting $A(e({\tt r}))=A(e({\tt w}))=A(e({\tt i}))=0$, and $A(e(0)=A(e(1))=1$, and setting $B(e(0))=B(e(1))=0$. Additional dimensions forward the current token to $h_{t}^{(1)}$. In the output of the first layer $z_{t}^{(1)}$, whenever the input is 0 or 1, the model now has access both to the current token $w_{t}$ and the preceding token $w_{t-1}$, which must have been an instruction. Based on this information, the model can set the gate to overwrite the state $h_{t-1}^{(2)}$ with the current input token when the preceding token was $\tt w$}, and pass along the state $h_{t-1}^{(2)}$ unaltered otherwise. This, together with $z_{t}^{(1)}$, is sufficient for always identifying the legal next symbols in $\mathcal{L}_{FF}$. The formal proof is in Appendix B.1.
Figure 2: (a) Visualizing the SSM equations 1, 2: The hidden state $H$ is updated by a combination of its previous values, transformed by matrix AA, and the input $X$, modulated by matrix $B$. The updated hidden state and input are then processed through a $Mix(.)$ layer, which can incorporate components like (Swi)GLU or Linear layers, with an optional RMSNorm for normalization. (b) An intuitive construction for recognizing PARITY with SSMs is achieved by setting $B=0$ and A=1A=-1 when the input is 11, and $A=1$ otherwise. However, this construction violates both nonnegative and time-invariant properties. We show that one of these properties is provably required to recognize PARITY at arbitrary lengths using an SSM (Theorem 2). (c) Modeling $a^{n}b^{n}$: the matrix AA adds the previous hidden state to the update, and depending on whether the input symbol requires counting up or down, matrix $B$ is set to 11 or 1-1, thus making the SSM simulate a counter (Theorem 5)